OpenStudy (anonymous):

Find the sales tax amount. purchase price = $78; tax rate = 6% A.$4.68 B.$82.68 C.$73.32 D.$82.32

OpenStudy (dan815):

how much is it with the tax

jimthompson5910 (jim_thompson5910):

Take 6% of 78 to get 6% of 78 = (6/100)*78 = ??

OpenStudy (dan815):

1% of 78$ = 0.78 cents so 6% = 0.78 * 6

OpenStudy (dan815):

thats how much money is added to your starting amount so calculate your total, good luck

OpenStudy (anonymous):

soo its a?

OpenStudy (yanasidlinskiy):

Divide just like the other ones said above. What is 6/100? After you get that multiply it by 78. And you get your answer!!:)